# Large Language Model Enhanced Recommender Systems: Taxonomy, Trend, Application and Future

### v2

@@ Abstract.
Recommender Systems; Large Language Model; Enhancement
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
The large language model has been proven to own unprecedented emergent capabilities in language understanding and reasoning (Zhao et al., 2023b; Chang et al., 2024; Wu et al., 2024a). In view of the only utilization of collaborative signals by conventional recommender systems (Bobadilla et al., 2013; Wang et al., 2022, 2024h), supplementing semantic information for RS by the LLM is attractive. Thus, many works have been proposed to fill the gap between natural language and recommendation, leading to a powerful RS.
@@
In order to clarify the range of this survey, we first give out the definition of LLMERS: The conventional recommender systems are enhanced by LLM via assistance in training or supplementary for data, while no need for LLM inference during the service. Though there have been some surveys about LLM for RS, three critical differences exist. i) Most current surveys focus on how to use the LLM itself as a better RS, including generative recommendation (Wu et al., 2024b; Li et al., 2024c, a) and discriminative recommendation (Lin et al., 2023; Zhao et al., 2024a; Chen et al., 2024c; Huang et al., 2024b; Vats et al., 2024; Chen et al., 2024b). By comparison, our survey addresses the LLMERS especially. ii) LLM for RS is a cutting-edge direction, which develops rapidly. Some surveys (Wu et al., 2024b; Lin et al., 2023; Zhao et al., 2024a; Chen et al., 2024b) do not include the most up-to-date papers. By comparison, more than $5050$ works in this survey were published after 2024. iii) A few surveys have referred to LLM-enhanced RS (Lin et al., 2023; Chen et al., 2024b), but they only focus on the enhancement from the aspect of feature engineering. In contrast, this survey is the first one to conclude LLMERS from a comprehensive view, including feature and model aspects.
Figure 1. The taxonomy of LLM-enhanced recommender systems.
@@ 1.1. Preliminary
Since LLM-enhanced RS is based on the conventional one, the introduction to its components and challenges is necessary to understand why and where the LLM is needed. As shown in Figure 1, the conventional RS often consists of the interaction data and recommendation model.
Interaction Data. The conventional recommender system captures the collaborative signals from user-item records (Koren et al., 2021), so the interactions in data are necessary for training. Besides, many content-based models (Lops et al., 2011) extract the co-occurrence relationships contained in the features of users and items for recommendation. Therefore, Feature and Interaction are two necessities of data. However, two challenges of data hinder further advancements for conventional RS.
@@ 1.2. Taxonomy
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.7cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2, , [Prompting Method, fill=gray!45, parent [Knowledge §2, for tree=pretrain [Summary Text §2.1, pretrain [User Only, pretrain [LANE (Zhao et al., 2024d); LLM-BRec (Jalan et al., 2024), pretrain_work] ] [Item Only, pretrain [ONCE (Liu et al., 2024a); LAMAR (Luo et al., 2024b); GPTAugNews (Yada and Yamana, 2024); SeRALM (Ren et al., 2024a); MMRec (Tian et al., 2024); X-Reflect (Lyu et al., 2024), pretrain_work] ] [User & Item, pretrain [KAR (Xi et al., 2024a); GaCLLM (Du et al., 2024); SLIM (Wang et al., 2024f); REKI (Xi et al., 2024b), pretrain_work] ] ] [Knowledge Graph §2.2, pretrain [Generation, pretrain [LLMRG (Wang et al., 2024a); SAGCN (Liu et al., 2023a); LLMHG (Chu et al., 2024); HRGraph (Wasi, 2024) , pretrain_work] ] [Completion & Fusion, pretrain [LLM-KERec (Zhao et al., 2024c); CSRec (Yang et al., 2024c), pretrain_work] ] ] [Combination §2.3, pretrain [KELLMRec (Luo et al., 2024a); SKarRec (Li et al., 2024b), pretrain_work] ] ] [Interaction §3, for tree=fill=red!45,template [Text-based §3.1, template [ONCE (Liu et al., 2024a); LLMRec (Wei et al., 2024); LlamaRec (Luo et al., 2024c); ColdAug (Wang et al., 2024e); LLMHD (Song et al., 2024), template_work] ] [Score-based §3.2, template [LLM-Ins (Huang et al., 2024a); LLM4DSR (Wang et al., 2024d), template_work] ] ] [Model
@@
Initialization §4.1, answer [Whole, answer [CTRL (Li et al., 2023); FLIP (Wang et al., 2024b); , answer_work] ] [Embedding, answer [LLM2X (Harte et al., 2023); SAID (Hu et al., 2024); LEARN (Jia et al., 2024); TSLRec (Liu et al., 2024g); LLMEmb (Liu et al., 2024d), answer_work] ] ] [Model
Distillation §4.2, answer [Feature, answer [LEADER (Liu et al., 2024f); SLMRec (Xu et al., 2024a), answer_work] ] [Response, answer [DLLM2Rec (Cui et al., 2024), answer_work] ] ] [Embedding Utilization §4.3, answer [User Only, answer [LLM-CF (Sun et al., 2024), answer_work] ] [Item Only, answer [TedRec (Xu et al., 2024b); LSVCR (Zheng et al., 2024); LRD (Yang et al., 2024b); NoteLLM (Zhang et al., 2024b); LLM-ESR (Liu et al., 2024e); CELA (Wang et al., 2024c); NoteLLM-2 (Zhang et al., 2024c);
AlphaRec (Sheng et al., 2024); Flare (Hebert et al., 2024); HLLM (Chen et al., 2024a), answer_work] ] [User & Item, answer [SINGLE (Liu et al., 2024b); LoID (Xu et al., 2024c); BAHE (Geng et al., 2024); DynLLM (Zhao et al., 2024b);
EmbSum (Zhang et al., 2024a); Laser (Lin et al., 2024); LARR (Wan et al., 2024), answer_work] ] ] [Embedding Guidance §4.4, answer [User Only, answer [LLM4SBR (Qiao et al., 2024); LLM-ESR (Liu et al., 2024e);
LLM4MSR (Wang et al., 2024g), answer_work] ] [User & Item, answer [RLMRec (Ren et al., 2024b); DaRec (Yang et al., 2024a), answer_work] ] ] ] ]
@@ 2.1.1. User Only and Item Only
For the user-only literature, LANE (Zhao et al., 2024d) and LLM-BRec (Jalan et al., 2024) fabricate prompts to capture multi-preferences of users. For better utilization, LANE additionally designs an attention mechanism to align the preference representations of RS and LLM. On the other hand, as basic units in the RS, more works focus on the item-only aspect. As one of the pioneers in this field, ONCE (Liu et al., 2024a) utilizes LLM as a content summarizer for news recommendation. Besides, it proposes to extend the news content by LLM before encoding, while Yada et al. (Yada and Yamana, 2024) uses the LLM to generate category descriptions for augmenting the news content. Similarly, LAMAR (Luo et al., 2024b) adopts LLM to output factual properties of movies. As for SeRALM (Ren et al., 2024a), it additionally devises an alignment training pattern for better integration of LLM’s knowledge. Recently, multimodal RS (Liu et al., 2024c) has attracted much attention. MMRec (Tian et al., 2024) and X-Reflect (Lyu et al., 2024) propose to use multimodal LLM to summarize the visual feature into textual description for use.
@@ 2.1.2. User & Item
Some research studies generate summary texts for both user and item to enhance RS, especially content-based RS (Javed et al., 2021). KAR (Xi et al., 2024a) is one of the pioneers in this line, which gets the reasoning knowledge of users and factual knowledge of items in textual form. The knowledge is then encoded by the designed multi-expert text encoders and integrated into the traditional CTR models, such as DeepFM (Guo et al., 2017) and DIN (Zhou et al., 2018). REKI (Xi et al., 2024b), a successor of KAR, further refines the knowledge extraction module for various scales of recommendation scenarios. Different from these two works, GaCLLM (Du et al., 2024) enhances the graph-based RS models by the knowledge of users and items. Specifically, it considers the LLM as an aggregator to extract structural information and thereby augment the RS.
@@ 2.3. Combination
In consideration of the effectiveness of these two types, some works also resort to combining both of them. KELLMRec (Luo et al., 2024a) is one representative, which integrates the KG for constructing prompts to avoid the hallucination problem when getting the summary texts from LLM. Besides, SKarRec (Li et al., 2024b) also links the entity in KG to the textual prompts, allowing the LLM to generate KG-aware knowledge that augments the concept recommendations.
@@ 3.1. Text-based
ONCE (Liu et al., 2024a) is the first to investigate the data augmentation for RS using LLM. It prompts the LLM to recommend news based on a user’s historical records. The generated news is added to the dataset as an augmented interaction for this user. However, the direct choice from a large item set is difficult for LLM. Thus, later works propose narrowing the selection to a smaller set of items for higher augmentation accuracy. For example, LLMRec (Wei et al., 2024) first adopts well-trained traditional RS to filter out a list of items and then prompt LLM to choose from them. Similarly, LlamaRec (Luo et al., 2024c) imposes the LLM to generate interactions from a random-sampled item set. Furthermore, Wang et al. (Wang et al., 2024e) proposes to utilize the LLM to judge which item a user prefers, generating pair-wise interactions for cold-start RS. In addition to generating new data, distinguishing noisy data is also a vital topic in data augmentation. LLMHD (Song et al., 2024) addresses this issue by applying the LLM to distinguish hard or noisy samples in sequential recommendation.
@@ 3.2. Score-based
As mentioned in the last subsection, it is difficult for LLM to select from amounts of items. Besides, the textual output often faces the out-of-corpus problem (Bao et al., 2023), i.e., the generated item name cannot correspond to anyone in the item set. To avoid these two problems, some works explore the score-based method. One typical work of this line is LLM-InS (Huang et al., 2024a). It firstly imposes the LLM to derive the semantic embeddings of users and items, and then generate augmentations by calculating the similarity between them. By comparison, LLMDSR (Wang et al., 2024d) considers the probability of word tokens as the confidence in deciding a noisy item and generates the new item by the similarity of embeddings.
@@ 4.1.1. Whole
CTRL (Li et al., 2023) and FLIP (Wang et al., 2024b) are the representatives in this line. CTRL first considers the semantic representation extracted from LLM and collaborative representation derived from the traditional RS model as two distinct modalities (Fang et al., 2022, 2023, 2024). It then aligns such two “modalities” by contrastive tasks. The parameters updated during the alignment process serve as the initialization of the whole traditional RS model. Building upon CTRL, FLIP (Wang et al., 2024b) fabricates fine-grained masked tabular and language modeling alignment tasks for better model initialization.
@@ 4.1.2. Embedding
Embedding plays a key role in the traditional RS models (Zhao et al., 2023a), preliminarily capturing the relationships between items and features. The LLMERS in this sub-category all aim to extract semantic relationships between items by the derived embeddings from LLM (i.e., LLM embeddings). Harte et al. (Harte et al., 2023) initiates this line, who input the item title to a large text embedding model. Then, the derived LLM embeddings are used to initialize the embedding layer of a sequential recommendation model. Later, LEARN (Jia et al., 2024) proposes to adopt the hidden states of a frozen open-sourced LLM as the item embedding. However, the general-purpose LLM without fine-tuning may encounter incompatible challenges for the recommendation task. To address this, SAID (Hu et al., 2024) and TSLRec (Liu et al., 2024g) specialize the item attribute word generation and information reconstruction tasks, respectively, to adapt the LLM to recommendation better. More recently, LLMEmb (Liu et al., 2024d) integrates an attribute-level contrastive task to further extract fine-grained semantic relationships between items.
@@ 4.2.2. Response-based
Following the advancement in KD for recommendation, recent works have explored the use of LLM to generate a ranking list from a semantic view, followed by the ranking distillation loss (Tang and Wang, 2018) for the RS model. Specifically, DLLM2Rec (Cui et al., 2024) proposes to fine-tune the LLM to adapt the recommendation task and then distill the small RS model by the ranking list derived from the LLM.
@@ 4.3.2. Item Only
In some types of recommender systems, e.g., sequential recommendation, the item embeddings play a fundamental role. Thus, most existing embedding utilization focuses on the item-only enhancement. Many early studies in this sub-category directly adopt the frozen LLM without fine-tuning. For example, TedRec (Xu et al., 2024b) apply LLaMA to encode the item texts and then propose a frequency-based fusion to combine the derived semantic embeddings and ID embeddings. LRD (Yang et al., 2024b) adopts LLM embeddings in a manner akin to TedRec but introduces a novel self-supervised relation discovery task to optimize the utilization. However, these works may lose the informative semantics contained in LLM embeddings during training. To address such an issue, LLM-ESR (Liu et al., 2024e) and AlphaRec (Sheng et al., 2024) both devise a new pattern of frozen LLM embeddings with a trainable adapter to maintain the original semantic information.
Despite their success in utilization, the general-purpose LLM is often not well suited for the recommendation tasks. Thus, several works propose to fine-tune the LLM before utilization. LEARN (Jia et al., 2024) proposes to align the user preferences derived from LLM with those of RS models, which refines the item LLM embeddings. NoteLLM (Zhang et al., 2024b) and NoteLLM-2 (Zhang et al., 2024c) are both devised for item-item recommendation, where the LLM derives the item embeddings for retrieval. Differently, NoteLLM only targets combining the textual semantics and collaborative signals, while NoteLLM-2 gives out a novel path to fine-tune the LLM for understanding the multimodal features of items. Recently, HLLM (Chen et al., 2024a) proposes a novel hierarchal framework for enhancement. The first level involves an item LLM, which encodes the textual features of the item. The derived item LLM embeddings then will be fed into a user LLM for final recommendation.

### v3

# Large Language Model Enhanced Recommender Systems: A Survey

@@ Abstract.
copyright: nonecopyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
The large language model has been proven to own unprecedented emergent capabilities in language understanding and reasoning (Zhao et al., 2023b; Chang et al., 2024; Wu et al., 2024a). In view of the only utilization of collaborative signals by conventional recommender systems (Bobadilla et al., 2013; Wang et al., 2022, 2024j), supplementing semantic information for RS by the LLM is attractive. Thus, many works have been proposed to fill the gap between natural language and recommendation, leading to a powerful RS.
@@
In order to clarify the range of this survey, we first give out the definition of LLMERS: The conventional recommender systems are enhanced by LLM via assistance in training or supplementary for data, while no need for LLM inference during the service. Though there have been some surveys about LLM for RS, three critical differences exist. i) Most current surveys focus on how to use the LLM itself as a better RS, including generative recommendation (Wu et al., 2024b; Li et al., 2024c, a) and discriminative recommendation (Lin et al., 2023; Zhao et al., 2024a; Chen et al., 2024d; Huang et al., 2024b; Vats et al., 2024; Chen et al., 2024b). By comparison, our survey addresses the LLMERS especially. ii) LLM for RS is a cutting-edge direction, which develops rapidly. Some surveys (Wu et al., 2024b; Lin et al., 2023; Zhao et al., 2024a; Chen et al., 2024b) do not include the most up-to-date papers. By comparison, more than $6060$ works in this survey were published after 2024. iii) A few surveys have referred to LLM-enhanced RS (Lin et al., 2023; Chen et al., 2024b), but they only focus on the enhancement from the aspect of feature engineering. In contrast, this survey is the first one to conclude LLMERS from a comprehensive view, including feature and model aspects.
@@ 1.1. Preliminary
Since LLM-enhanced RS targets to enhance RS, introducing RS’s components and challenges is a prerequisite to understanding where and why the LLM is needed. As shown in Figure 1, the conventional RS often consists of the interaction data (feature and interaction) and recommendation model.
Interaction Data. The conventional recommender system captures the collaborative signals from user-item records (Koren et al., 2021), which depends on the interactions in data for training. Besides, many content-based models (Lops et al., 2011) extract the co-occurrence relationships contained in the features of users and items for recommendation. Therefore, Feature and Interaction are two necessities of data. However, there exist two challenges of data hindering further advancements for conventional RS.
@@ 1.2. Taxonomy
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.7cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2, , [Prompting Method, fill=gray!45, parent [Knowledge §2, for tree=pretrain [Summary Text §2.1, pretrain [User Only, pretrain [LANE (Zhao et al., 2024d); LLM-BRec (Jalan et al., 2024); RecLM (Jiang et al., 2024) , pretrain_work] ] [Item Only, pretrain [ONCE (Liu et al., 2024a); LAMAR (Luo et al., 2024d); GPTAugNews (Yada and Yamana, 2024); SeRALM (Ren et al., 2024a); MMRec (Tian et al., 2024); X-Reflect (Lyu et al., 2024), pretrain_work] ] [User & Item, pretrain [KAR (Xi et al., 2024a); GaCLLM (Du et al., 2024); SLIM (Wang et al., 2024h); REKI (Xi et al., 2024b), pretrain_work] ] ] [Knowledge Graph §2.2, pretrain [Generation, pretrain [LLMRG (Wang et al., 2024a); SAGCN (Liu et al., 2023a); LLMHG (Chu et al., 2024); HRGraph (Wasi, 2024); LLM-PKG (Wang et al., 2024b); AutoGraph (Shan et al., 2024); CIKG (Hu et al., 2024a) , pretrain_work] ] [Completion & Fusion, pretrain [LLM-KERec (Zhao et al., 2024c); CSRec (Yang et al., 2024c); CoLaKG (Cui et al., 2024b), pretrain_work] ] ] [Combination §2.3, pretrain [KELLMRec (Luo et al., 2024c); SKarRec (Li et al., 2024b), pretrain_work] ] ] [Interaction §3, for tree=fill=red!45,template [Text-based §3.1, template [ONCE (Liu et al., 2024a); LLMRec (Wei et al., 2024); LlamaRec (Luo et al., 2024e); ColdAug (Wang et al., 2024f); LLMHD (Song et al., 2024a); LLM4IDRec (Chen et al., 2024c); SampleLLM (Gao et al., 2025) , template_work] ] [Score-based §3.2, template [LLM-Ins (Huang et al., 2024a); LLM4DSR (Wang et al., 2024e); EIMF (Qiao et al., 2024a) , template_work] ] ] [Model
@@
Initialization §4.1, answer [Whole, answer [CTRL (Li et al., 2023); FLIP (Wang et al., 2024c); , answer_work] ] [Embedding, answer [LLM2X (Harte et al., 2023); SAID (Hu et al., 2024b); LEARN (Jia et al., 2024); TSLRec (Liu et al., 2024g); LLMEmb (Liu et al., 2024d); LLMInit (Zhang et al., 2025) , answer_work] ] ] [Model
Distillation §4.2, answer [Feature, answer [LEADER (Liu et al., 2024f); SLMRec (Xu et al., 2024a), answer_work] ] [Response, answer [DLLM2Rec (Cui et al., 2024a), answer_work] ] ] [Embedding Utilization §4.3, answer [User Only, answer [LLM-CF (Sun et al., 2024); LLMGPR (Long et al., 2024) , answer_work] ] [Item Only, answer [TedRec (Xu et al., 2024b); LSVCR (Zheng et al., 2024); LRD (Yang et al., 2024b); NoteLLM (Zhang et al., 2024b); LLM-ESR (Liu et al., 2024e); CELA (Wang et al., 2024d); NoteLLM-2 (Zhang et al., 2024c);
AlphaRec (Sheng et al., 2024); Flare (Hebert et al., 2024); HLLM (Chen et al., 2024a); QARM (Luo et al., 2024a); Molar (Luo et al., 2024b); PRECISE (Song et al., 2024b); PAD (Wang et al., 2024g); RecG (Li et al., 2025) , answer_work] ] [User & Item, answer [SINGLE (Liu et al., 2024b); LoID (Xu et al., 2024c); BAHE (Geng et al., 2024); DynLLM (Zhao et al., 2024b);
EmbSum (Zhang et al., 2024a); Laser (Lin et al., 2024); LARR (Wan et al., 2024), answer_work] ] ] [Embedding Guidance §4.4, answer [User Only, answer [LLM4SBR (Qiao et al., 2024b); LLM-ESR (Liu et al., 2024e);
LLM4MSR (Wang et al., 2024i), answer_work] ] [User & Item, answer [RLMRec (Ren et al., 2024b); DaRec (Yang et al., 2024a); HARec (Ma et al., 2024); IRLLRec (Wang et al., 2025) , answer_work] ] ] ] ]
@@ 2.1.1. User Only and Item Only
For the user-only literature, LANE (Zhao et al., 2024d) and LLM-BRec (Jalan et al., 2024) fabricate prompts to capture multi-preferences of users. For better utilization, LANE additionally designs an attention mechanism to align the preference representations of RS and LLM. On the other hand, as basic units in the RS, more works focus on the item-only aspect. As one of the pioneers in this field, ONCE (Liu et al., 2024a) utilizes LLM as a content summarizer for news recommendation. Besides, it proposes to extend the news content by LLM before encoding, while Yada et al. (Yada and Yamana, 2024) uses the LLM to generate category descriptions for augmenting the news content. Similarly, LAMAR (Luo et al., 2024d) adopts LLM to output factual properties of movies. As for SeRALM (Ren et al., 2024a), it additionally devises an alignment training pattern for better integration of LLM’s knowledge. Recently, multimodal RS (Liu et al., 2024c) has attracted much attention. MMRec (Tian et al., 2024) and X-Reflect (Lyu et al., 2024) propose to use multimodal LLM to summarize the visual feature into textual description for use.
@@ 2.1.2. User & Item
Some studies generate summary texts for both user and item to enhance RS, especially content-based RS (Javed et al., 2021). As one of the pioneers in this line, KAR (Xi et al., 2024a) gets the reasoning knowledge of users and factual knowledge of items in textual form. The knowledge is then encoded by the designed multi-expert text encoders and integrated into the traditional CTR models, such as DeepFM (Guo et al., 2017) and DIN (Zhou et al., 2018). REKI (Xi et al., 2024b), a successor of KAR, further refines the knowledge extraction module for various scales of recommendation scenarios. Different from these two works, GaCLLM (Du et al., 2024) enhances the graph-based RS models by the knowledge of users and items, considering the LLM as an aggregator to extract structural information and thereby augment the RS.
@@ 2.3. Combination
In consideration of the effectiveness of these two types, some works also resort to combining both of them. KELLMRec (Luo et al., 2024c) is one representative, which integrates the KG for constructing prompts to avoid the hallucination problem when getting the summary texts from LLM. Besides, SKarRec (Li et al., 2024b) also links the entity in KG to the textual prompts, allowing the LLM to generate KG-aware knowledge that augments the concept recommendations.
@@ 3.1. Text-based
ONCE (Liu et al., 2024a) is the first to investigate the data augmentation for RS using LLM. It prompts the LLM to recommend news based on a user’s historical records. The generated news is added to the dataset as an augmented interaction for this user. However, the direct choice from a large item set is difficult for LLM. Thus, later works propose narrowing the selection to a smaller set of items for higher augmentation accuracy. For example, LLMRec (Wei et al., 2024) first adopts well-trained traditional RS to filter out a list of items and then prompt LLM to choose from them. Similarly, LlamaRec (Luo et al., 2024e) imposes the LLM to generate interactions from a random-sampled item set. Furthermore, Wang et al. (Wang et al., 2024f) proposes to utilize the LLM to judge which item a user prefers, generating pair-wise interactions for cold-start RS. In addition to generating new data, distinguishing noisy data is also a vital topic in data augmentation. LLMHD (Song et al., 2024a) addresses this issue by applying the LLM to distinguish hard or noisy samples in sequential recommendation.
@@ 3.2. Score-based
As mentioned in the last subsection, it is difficult for LLM to select from amounts of items. Besides, the textual output often faces the out-of-corpus problem (Bao et al., 2023), i.e., the generated item name cannot correspond to anyone in the item set. To avoid these two problems, some works explore the score-based method. One typical work of this line is LLM-InS (Huang et al., 2024a). It firstly imposes the LLM to derive the semantic embeddings of users and items, and then generate augmentations by calculating the similarity between them. By comparison, LLMDSR (Wang et al., 2024e) considers the probability of word tokens as the confidence in deciding a noisy item and generates the new item by the similarity of embeddings.
@@ 4.1.1. Whole
CTRL (Li et al., 2023) and FLIP (Wang et al., 2024c) are the representatives in this line. CTRL first considers the semantic representation extracted from LLM and collaborative representation derived from the traditional RS model as two distinct modalities (Fang et al., 2022, 2023, 2024). It then aligns such two “modalities” by contrastive tasks. The parameters updated during the alignment process serve as the initialization of the whole traditional RS model. Building upon CTRL, FLIP (Wang et al., 2024c) fabricates fine-grained masked tabular and language modeling alignment tasks for better model initialization.
@@ 4.1.2. Embedding
Embedding plays a key role in the traditional RS models (Zhao et al., 2023a), preliminarily capturing the relationships between items and features. The LLMERS in this sub-category all aim to extract semantic relationships between items by the derived embeddings from LLM (i.e., LLM embeddings). Harte et al. (Harte et al., 2023) initiates this line, who input the item title to a large text embedding model. Then, the derived LLM embeddings are used to initialize the embedding layer of a sequential recommendation model. Later, LEARN (Jia et al., 2024) proposes to adopt the hidden states of a frozen open-sourced LLM as the item embedding. However, the general-purpose LLM without fine-tuning may encounter incompatible challenges for the recommendation task. To address this, SAID (Hu et al., 2024b) and TSLRec (Liu et al., 2024g) specialize the item attribute word generation and information reconstruction tasks, respectively, to adapt the LLM to recommendation better. More recently, LLMEmb (Liu et al., 2024d) integrates an attribute-level contrastive task to further extract fine-grained semantic relationships between items.
@@ 4.2.2. Response-based
Following the advancement in KD for recommendation, recent works attempt to use LLM to generate a ranking list from a semantic view, followed by the ranking distillation loss (Tang and Wang, 2018) for the RS model. Specifically, DLLM2Rec (Cui et al., 2024a) proposes to fine-tune the LLM to adapt the recommendation task and then distill the small RS model by the ranking list derived from the LLM.
@@ 4.3.2. Item Only
In some types of recommender systems, e.g., sequential recommendation, the item embeddings play a fundamental role. Thus, most existing embedding utilization focuses on the item-only enhancement. Many early studies in this sub-category directly adopt the frozen LLM without fine-tuning. For example, TedRec (Xu et al., 2024b) applies LLaMA to encode the item texts and then proposes a frequency-based fusion to combine the derived semantic embeddings and ID embeddings. LRD (Yang et al., 2024b) adopts LLM embeddings akin to TedRec but introduces a novel self-supervised relation discovery task to optimize the utilization. However, these works may lose the informative semantics of LLM embeddings during training. To address such issue, LLM-ESR (Liu et al., 2024e) and AlphaRec (Sheng et al., 2024) both devise a new pattern of frozen LLM embeddings with a trainable adapter to maintain the original semantic information.
Despite their successful utilization, the general-purpose LLM is often not well suited for the recommendation tasks. Thus, several works propose to fine-tune the LLM before utilization. LEARN (Jia et al., 2024) proposes to align the user preferences derived from LLM with those of RS models, which refines the item LLM embeddings. NoteLLM (Zhang et al., 2024b) and NoteLLM-2 (Zhang et al., 2024c) are both devised for item-item recommendation, where the LLM derives the item embeddings for retrieval. NoteLLM only targets combining the textual semantics and collaborative signals, while NoteLLM-2 gives out a novel path to fine-tune the LLM for understanding the multimodal features of items. Recently, HLLM (Chen et al., 2024a) proposes a novel hierarchal framework for enhancement. The first level involves an item LLM, which encodes the textual features of the item. The item LLM embeddings will then be fed into a user LLM for final recommendation.